OpenStudy (anonymous):

GIVING MEDAL N FAN determine if the graph is symmetric about the x-axis, the y-axis or the orgiin r= 5 cos 5theta y-axis only x-axis, y-axis x-axis, y-axis, origin x-axis only

OpenStudy (anonymous):

@Michele_Laino

OpenStudy (welshfella):

do you know what the graph of cos x looks like?

OpenStudy (anonymous):

NO

OpenStudy (welshfella):

if f(x) = f(-x) then its symetrical about the y-axis

OpenStudy (michele_laino):

using the De Moivre formula (complex numbers), we get this identity: \[\large \cos \left( {5\theta } \right) = {\left( {\cos \theta } \right)^5} - 10{\left( {\cos \theta } \right)^3}{\left( {\sin \theta } \right)^2} + 5\left( {\cos \theta } \right){\left( {\sin \theta } \right)^4}\]

OpenStudy (anonymous):

y axis only @Michele_Laino ?

OpenStudy (michele_laino):

so, we can rewrite your equation as follows: \[\large r = \left\{ {5{{\left( {\cos \theta } \right)}^5} - 10{{\left( {\cos \theta } \right)}^3}{{\left( {\sin \theta } \right)}^2} + 5\left( {\cos \theta } \right){{\left( {\sin \theta } \right)}^4}} \right\}\]

OpenStudy (michele_laino):

and going to the cartesian coordinates, we can write: \[\Large r = \left\{ {5\frac{{{x^5}}}{{{r^5}}} - 10\frac{{{x^3}{y^2}}}{{{r^5}}} + 5\frac{{x{y^4}}}{{{r^5}}}} \right\}\] so what can you conclude?

OpenStudy (anonymous):

is it all 3? x-axis, y-axis, origin

OpenStudy (michele_laino):

oops... I have made a typo, here is the right equation: \[\Large r = 5\left\{ {\frac{{{x^5}}}{{{r^5}}} - 10\frac{{{x^3}{y^2}}}{{{r^5}}} + 5\frac{{x{y^4}}}{{{r^5}}}} \right\}\] we have a symmetry for a variable whose exponent is an even number

OpenStudy (anonymous):

omg im lost is it x-axis or y-axis, could u just tell me ? @Michele_Laino

OpenStudy (welshfella):

check if f(-60) = f(60) if so then its symmetrical over the y-axis

OpenStudy (michele_laino):

I think x-axis, since we have this: \[\Large f\left( {x, - y} \right) = f\left( {x,y} \right)\]

OpenStudy (anonymous):

so x-axis everyone @Michele_Laino @welshfella @Astrophysics ??

OpenStudy (welshfella):

no i think its y-axis

OpenStudy (anonymous):

both??

OpenStudy (anonymous):

or just y axis?

OpenStudy (welshfella):

just y-axis

OpenStudy (anonymous):

ok

OpenStudy (welshfella):

because its an even function and graph of cos x is